OpenStudy (anonymous):

Please Solve: log(base c) of 81=4

myininaya (myininaya):

\[c^4=81\]

myininaya (myininaya):

ln(c^4)=ln(81) 4ln(c)=ln(81)

myininaya (myininaya):

divide 4 on both sides then ln(c)=ln(81)/4 e^ e^

myininaya (myininaya):

i cant see some of my screen my cat is in front of it

OpenStudy (anonymous):

*cough* 3^4=81 *uncough*

myininaya (myininaya):

shh.... my cat is messing with my concentration

OpenStudy (anonymous):

lolol i had to kick the cat out of my room for a bit so i could get work done.

myininaya (myininaya):

by the way blooz if you didn't pick up what joe was throwing down he was saying skip all that nonsense because c is simply 4 since 3^4=81

OpenStudy (karatechopper):

i dont think myin is tht harsh to do tht joe...... jk!:)